Micromedia offers computer training seminars on a variety of topics. In the seminars each student works at a personal computer, practicing the particular activity that the instructor is presenting. Micromedia is currently planning a two-day seminar on the use of Microsoft Excel in statistical analysis. The projected fee for the seminar is $600 per student. The cost for the conference room, instructor compensation, lab assistants, and promotion is $9600 per 2 days. Micromedia rents computers for its seminars at a cost of $120 per computer per day. Develop a model for the total cost to put on the seminar. Let x represent the number of students who enroll in the seminar. Round your answers to the nearest whole number. For subtractive or negative number use a minus sign even if there is a + sign before the blank. (Example: -300) Total cost = 4,800 + ·x Develop a model for the total profit if x students enroll in the seminar. Round your answers to the nearest whole number. For subtractive or negative number use a minus sign even if there is a + sign before the blank. (Example: -300) Total profit = 240 ·x + -4,800 Micromedia has forecasted an enrollment of 30 students for the seminar. How much profit will be earned if their forecast is accurate? Round your answer to the nearest dollar. Total profit = $ 1,200 Compute the breakeven point. Round your answer to the nearest whole number. The breakeven point is students.
Answer #1

Breakdown point is point where input cost becomes equal to that of output, profit is zero at breakdown point.

Total cost for 2 days = 9600 + 2* 120x = 9600 + 240x

Total output from seminar = 600 * x

When input = output,

9600 + 240x = 600x

Solving for x: x = 26.67

Rounding to nearest whole number 27

If forcasted number of participants is 30,

Profit = 600*30 -(9600+ 240*30) = 1200 dollars

Now,

If total profit is 1200, breakdown point calculation

1200= 600x - (9600 + 240x)

Solving for x : x = 30 participants.